State how many imaginary and real zeros the function has.

f(x) = x4 - 15x2 - 16

4 imaginary; 0 real

3 imaginary; 1 real

2 imaginary; 2 real

0 imaginary; 4 real


Help me please and show work
1
Expert's answer
2014-06-03T05:46:01-0400
Assume that y=x^2. We obtain y^2-15y-16=(y-16)(y+1). So, x^2=16 and x^2=-1.

Thus, zeros are 4,-4,i and -i.

And the answer is 2 imaginary; 2 real.